Jackie is five years older than her brother. In four years, Jackie will be twice as old as her brother will be. How old is Jackie now?
Answer by ptaylor(2198)

Let x=Jackie's brother's age now; in 4 years he will be x+4 years old
Then x+5=Jackie's age now; in 4 years she will be x+5+4=x+9 years old
Now we are told that:
x+9=2(x+4) get rid of parens (distributive)
x+9=2x+8 subtract x and also 8 from each side
x-x+9-8=2x-x+8-8 collect like terms
1=x or
x=1---------------------Jackie's brother's age now
x+5=1+5=6----------------Jackie's age now
CK
6+4=2*(1+4)
10=10
Hope this helps----ptaylor